OpenStudy (anonymous):

2sqrt2costheata+4=6 Is this pi/4 and 7pi/4

OpenStudy (anonymous):

\[2\sqrt{2}\cos \theta+4=6\]

OpenStudy (michele_laino):

we have to subtract 4 from both sides first

OpenStudy (anonymous):

I did that then I divided by 4

OpenStudy (michele_laino):

no, please after that subtraction, you have to divide by 2*sqrt(2)

OpenStudy (anonymous):

I meant 2 my bad

OpenStudy (michele_laino):

now divide by sqrt(2)

OpenStudy (michele_laino):

you should get this: \[\cos \theta = \frac{1}{{\sqrt 2 }}\]

OpenStudy (anonymous):

Is this answer then pi/4 and 7pi/4?

OpenStudy (michele_laino):

ok!

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

actually need to redo that one.. one sec

OpenStudy (jdoe0001):

have a few typos darn it

OpenStudy (jdoe0001):

\(\bf 2\sqrt{2}cos (\theta)+4=6\implies 2\sqrt{2}cos (\theta)=2 \\ \quad \\ cos (\theta)=\cfrac{\cancel{2}}{\cancel{2}\sqrt{2}}\implies cos^{-1}[cos(\theta)]=cos^{-1}\left( \cfrac{1}{\sqrt{2}} \right) \\ \quad \\ \theta=cos^{-1}\left( \cfrac{1}{\sqrt{2}} \right)\)

OpenStudy (michele_laino):

ok! your solution is right for the first round angle @minisweet4